Journalist Report – December 22nd

[Sol 4]

And we are back to red! After 24h of snow and landscapes looking more like Europa than Mars, the sun is shinning again and the temperatures are positive.

Today, we did a short walking EVA to pick up geological samples around the Hab and to see if X-1 is capable of doing it. We brought back a dozen interesting rocks and samples of soil.

Then, this afternoon, we practiced CPR with John to be sure that we would be able to do it if needed. We all managed to rescue Phil the mannequin successfully!

The crew is doing great so far. The only concern is that our water supplies are running low, even though we are careful and didn’t have a shower yet. We discussed tonight how to make it even lower. Hopefully, we will have enough water for the rest of the mission. Think of us when you will have your next shower, it is a dream here!

Greetings from the Red Planet,

Thibault, ExO and spacesuit engineer for the Crew 185

Greenhab Report – December 22nd

Greenhab Report
David Murray
22 Dec, 2017
Environmental control: (Choose which is appropriate and explain further if needed)
Ambient with door open from about noon to 15:30
Working Hour: 15:30
Inside temp at working hour: 24 C
Outside temp during working hours: -0.5 C
Inside temperature high: 33 C
Inside temperature low: 14 C
Inside humidity: 16%
Functioning nominally,
Turns on at 16 C and off at 20 C
Not functional
Shade Cloth On/Off
The greenhab lights are on from 17:00 to 24:00 every night
Average temperature:
Changes to Crops
Morning Research Observations

Daily Water Usage for Crops
4 gallons + 3 gallons of miracle grow and water
Time(s) of Watering Crops
I transplanted some more tomatoes and applied Miracle Grow fertilizer to all the plants in the greenhab with the exception of my control and experimental groups of tomato plants. I added tomato cages to the bean plants for some support because that’s the best option I have right now. All the plants are looking healthy.
Support/Supplies Needed

Operations Report – December 22nd

Crew 185

Operations Report Friday 12/22/2017

SOL: 4

Name of person filing report: A.Passeron

Non-nominal systems:

-Spacesuit N°3 :
The ventilation system worked properly during the whole EVA. The running time of the fully charged battery is thus more than 1h20min.

– Shower-head :
Shower-head leak was fixed at 9:10am.

– Water shortage :
Crew aware of concern about water shortage. We are now using jerrycans for drinking water.

Generator (hours run):
Generator turned off at 8:10 am (SOC at 100%)
Generator turned on at 5:00pm (SOC at 90%)

Solar— SOC
@ 8:10am : 100%
@ 9:06am : 95%
@ 10:00am : 100%
@ 11:00am : did not check because of EVA
@ 12:00pm : 100%
@ 1:05pm : 100%
@ 2:35pm : 100%
@ 3:30pm : 99%
@ 5:00pm : 90%
@ 5:40pm : 90%

Diesel – 80%

Propane – 54%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 3.5 Gallons

Water (trailer) – 50 gallons

Water (static) – 280 gallons

Trailer to Static Pump used – No

Static to Loft Pump used – Yes

Water Meter: 50 gallons

Toilet tank emptied: No

ATVs Used: No
Oil Added? No
ATV Fuel Used:
# Hours the ATVs were used today: 00:00 hours
Notes on ATVs: ATVs were nominal.

Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes

Opportunity rover used: No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes

Curiosity rover used: No
Beginning charge:
Ending charge:
Currently charging:

HabCar used and why, where? No

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: Shower-head leak is fixed. Water shortage prevention.

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support:

Copyright © The Mars Society. All rights reserved. | Main Site